OpenStudy (anonymous):

Solve the system y = x + 2 and y = –x – 3 by using graph paper or graphing technology.

sam (.sam.):

y = x + 2 and y = –x – 3 Equate each other, x + 2= –x – 3 2x+2=-3 2x=-5 x=-(5)/(2) -------------------------------- Simply pick one equation, y = x + 2 y = -(5)/(2) + 2 y=-(1)/(2)

OpenStudy (anonymous):

thank you